A hawk is hunting and sees a mouse 100 m away. The hawk catches the mouse in
12.2s. The speed of the hawk was recorded at 35.8 m/s. What was the rate of
acceleration?
1
Expert's answer
2020-05-27T10:18:03-0400
"a=\frac{v}{t}"

"a=\frac{35.8}{12.2}=2.93\frac{m}{s^2}"